# Establishing validated standards for Home and Work location Detection

### v1

@@ Abstract
Smartphone location data have transformed urban mobility research, providing unprecedented insights into how people navigate and interact in cities. However, leveraging location data at scale presents methodological challenges. Accurately identifying individuals’ home and work locations is critical for a range of applications, including commuting analysis, unemployment estimation, and urban accessibility studies. Despite their widespread use, home–work detection methods lack a standardized framework that accounts for differing data quality and that is validated against ground-truth observations. This limits the comparability and reproducibility of results across studies and datasets. In this paper, we present HoWDe, a robust algorithm for identifying home and work locations from mobility data, explicitly designed to handle missing data and varying data quality across individuals. Using two unique ground-truth datasets comprising over $5,100$ individuals from more than $80$ countries, HoWDe achieves home and work detection accuracies of up to $97\%$ and $88\%$, respectively, with consistent performance across countries and demographic groups. We examine how parameter choices shape the trade-off between accuracy and user retention, and demonstrate how these methodological decisions influence downstream applications such as employment estimation and commuting pattern analysis. By supporting in-house pre-processing through a transparent and validated pipeline, HoWDe also facilitates the sharing of privacy-preserving mobility data. Together, our tools and findings establish methodological standards that support more robust, scalable, and reproducible mobility research at both individual and urban scales.
Keywords: GPS data, Home and work detection, Human Mobility, Stop-location sequences, Mobility algorithms, Missing data, Urban analytics
@@ 1 Introduction
Location data collected from smartphones has become indispensable for human mobility analyses [1, 2]. These datasets allow studying large samples of individuals globally, providing insights on fundamental aspects of human movements [3], epidemic spreading [4, 5, 6, 7], migration patterns [8, 5], transportation modeling [9], urban planning [10, 11], and the impact of natural disasters and societal shocks on population movements [12, 13, 14, 15, 16]. In light of these advancements, developing robust methods for location data analysis has become a pressing need for scientists and practitioners [17].
@@
A critical component of location data analyses involves the detection of individuals’ visits to their residential [18, 19, 20], and work location [17]. The accurate identification of home and work locations is key for mobility research, as it enables the analysis of factors such as socio-economic background [21, 22, 15], commuting behaviors [23, 16], social contact networks [24, 14], and accessibility to amenities and services [25, 14]. More broadly, it contributes to advancements in computational demography and social science research [26, 17].
@@
The most widely used methods for home and work detection are heuristic approaches that leverage the regularity of human behavior, such as the tendency to spend nights at home and working hours at a workplace [27, 28, 29]. These methods face significant challenges. First, they are sensitive to missing data, which is common in smartphone records [17]. Data quality varies across individuals and over time, as location gaps arise due to factors like inactivity, disabled services, or poor signal, introducing biases in home and work inference [30, 31, 17]. Second, they rely on the choice of parameters, such as time span required to establish a home location or the minimum number of hours per day spent at work.
@@
Despite their importance for mobility analyses, the extent to which heuristic approaches for home/work location detection are reliable remains unclear, and there is no consensus on how to choose one specific algorithm and its parameters’ values [21, 32, 33, 34, 35, 17], in part due to a lack of ground truth data. Validation efforts so far have been limited, focusing on small subsets of individuals, home detection alone, and datasets with low spatial resolution and short time periods [34, 35, 32, 21]. Furthermore, the influence of key parameter choices (such as the length of the period considered or the minimum activity level necessary to define home/work locations) has not been understood in relation to algorithmic performance. The lack of thorough validation has hindered the refinement and harmonization of these methods into standardized best practices [36, 17]. In turn, the absence of such standards can impact the comparability and reproducibility of mobility studies [17].
@@
In this work, we leverage two distinct ground truth datasets to address these key challenges in home/work location detection. Both datasets include sequences of visits to detailed locations with corresponding ground truth: the first dataset (D1) contains self-reported home and work locations from $4,599$ users collected between 2015 and 2019, while the second (D2) is based on independent expert annotations for $500$ users throughout 2020 (see Methods for details).
First, we leverage these data to uncover behavioral patterns in home and work visits. This allows us to empirically assess whether the assumptions underlying home and work detection hold in practice. Additionally, it enables us to quantify the variability in home-work visit behavior across a diverse set of users.
@@
Second, we formalize HoWDe, an algorithm for home and work detection that enhances existing heuristics by integrating key elements from previous work into a comprehensive approach. HoWDe combines the explicit detection of both home and work locations [14, 29], the identification of the workplace based on recurrent visits during working hours [29], the identification of the home as the most frequently visited location at night [35, 28, 34, 21], a sliding window to identify home and job changes over time [14], with tunable window length [14]. Additionally, HoWDe improves upon existing approaches by better accounting for data gaps and their variation across users and over time. We provide an open-source implementation of the algorithm [37].

### v2

# HoWDe: a validated algorithm for Home and Work location Detection

@@ Abstract
Smartphone location data have become a key resource for understanding urban mobility, yet extracting actionable insights requires robust and reproducible preprocessing pipelines. A central step is the identification of individuals’ home and work locations, which underpins analyses of commuting, employment, accessibility, and socioeconomic patterns. However, existing approaches are often ad hoc, data-specific, and difficult to reproduce, limiting comparability across studies and datasets. We introduce HoWDe, an open-source software library for detecting home and work locations from large-scale mobility data. HoWDe implements a transparent, modular pipeline explicitly designed to handle missing data, heterogeneous sampling rates, and differences in data sparsity across individuals. The code allows users to tune a small set of interpretable parameters, enabling to adapt the algorithm to diverse applications and datasets. Using two unique ground truth datasets comprising 5,0995{,}099 individuals across $68$ countries, we show that HoWDe achieves home and work detection accuracies of up to $97\%$ and $88\%$, respectively, with consistent performance across demographic groups and geographic contexts. We further demonstrate how parameter settings propagate to downstream metrics such as employment estimates and commuting flows, highlighting the importance of transparent methodological choices. By providing a validated, documented, and easily deployable pipeline, HoWDe supports scalable in-house preprocessing and facilitates the sharing of privacy-preserving mobility datasets. Our software and evaluation benchmarks establish methodological standards that enhance the robustness and reproducibility of human mobility research at urban and national scales.
Keywords: Home and work detection, Open-source, Human Mobility, GPS data, Stop-location data, Missing data, Urban analytics
@@ 1 Introduction
Location data collected from smartphones has become a key resource for the analysis of urban systems and human mobility [1, 2, 3, 4, 5, 6]. These datasets enable the study of large, diverse populations and have been used within a range of areas, including the study of human movements [7, 8, 9], epidemic spreading [10, 11, 12, 13], migration patterns [14, 11], transportation modeling [15], urban planning [16, 4, 5, 6], and assessing the impact of natural disasters and societal shocks on population movements [17, 18, 19, 20, 8]. As the availability of such data increases, developing robust methods for location data analysis has become a pressing need for scientists and practitioners [21].
@@
A critical component of location data analyses is the detection of individuals’ residential [22, 23, 24], and work locations [21]. The accurate identification of home and work locations is key for mobility research, as it enables the analysis of factors such as socio-economic background [25, 26, 20], commuting behaviors [27, 8], social contact networks [28, 19], and accessibility to amenities and services [29, 19]. More broadly, it contributes to advancements in computational demography and social science research [30, 21].
@@
The most widely used methods for home and work detection are heuristic approaches that leverage the regularity of human behaviour, such as the tendency to spend nights at home and working hours at a workplace [31, 32, 33]. However, these methods face significant challenges. First, they are sensitive to missing data, which often stems from inactivity or poor signal and can bias home and work detection [34, 35, 21]. Second, they rely on the choice of parameters, such as the time span required to establish a home location or the minimum number of visits required to identify a valid work location. These limitations underscore the need for methods that systematically address both the heterogeneity of missing data within individuals and the sensitivity to heuristic parameter definitions [36, 21].
@@
Despite their importance for mobility analyses, the extent to which heuristic approaches for home/work location detection are reliable remains insufficiently validated, and there is no consensus on how to choose one specific algorithm and its parameters’ values [25, 37, 38, 39, 40, 21], in part due to a scarcity of ground truth data. Existing validation efforts have been limited to small samples, short time periods, low spatial resolution, or to home detection alone [39, 41, 40, 37, 25]. Furthermore, the influence of key parameter choices (such as the length of the period considered or the minimum activity level necessary to define home/work locations) has not been studied systematically. Limited validation has hindered the refinement and harmonisation of these methods into standardised best practices  [21]. In turn, the absence of standard algorithms hinders the comparability and reproducibility of mobility studies  [42, 21].
@@
In this study, we leverage two distinct ground truth datasets collected across more than 68 countries to address these key challenges in home/work location detection. Both datasets include sequences of visits to detailed locations with corresponding ground truth: the first dataset (D1) contains self-reported home and work locations from $4,599$ users collected between 2015 and 2019, while the second (D2) is based on independent expert annotations for $500$ users throughout 2020 and 2021.
@@
First, we formalise HoWDe, an algorithm for home and work detection that enhances existing heuristics by integrating key elements from previous work into a comprehensive approach. HoWDe combines the explicit detection of both home and work locations [19, 33], the identification of the workplace based on recurrent visits during working hours [33], the identification of the home as the most frequently visited location at night [40, 32, 39, 25], a sliding window to identify home and job changes over time [19], with tunable window length [19]. Additionally, HoWDe improves upon existing approaches by accommodating users whose data sparsity varies over time. This is achieved by identifying home and work locations based on the proportion of time spent at a given location, while controlling for individual-level sparsity rather than excluding users based on absolute time or visit thresholds, as in prior work.